Write down a C program to find the solution of following system of equations using Gauss-Jordan Method and correct up to ten decimal points

5x-3y +72 = 4
3x + 26 y + 2z = 9
7x+2y +8z = 4
1
Expert's answer
2021-10-13T04:57:20-0400
5     3     7     4
 3    26     2     9
 7     2    10     5
Multiply row 1 by 1/5:
 1   3/5   7/5   4/5
 3    26     2     9
 7     2    10     5
Add -3 times row 1 to row 2:
 1    3/5   7/5   4/5
 0  121/5 -11/5  33/5
 7     2    10     5
Add -7 times row 1 to row 3:
 1    3/5   7/5   4/5
 0  121/5 -11/5  33/5
 0  -11/5   1/5  -3/5
Multiply row 2 by 5/121:
 1    3/5   7/5   4/5
 0      1 -1/11  3/11
 0  -11/5   1/5  -3/5
Add -3/5 times row 2 to row 1:
 1      0  16/11  7/11
 0      1  -1/11  3/11
 0  -11/5   1/5  -3/5
Add 11/5 times row 2 to row 3:
 1      0  16/11  7/11
 0      1  -1/11  3/11
 0      0    0     0

This is as far as we can go. This translates into:

x + (16/11)z = 7/11 or x = (-16/11)z + 7/11